prove that sin2thetha/1+cos2thetha=tanthetha

Answer by Alan3354(69443)

prove that sin2thetha/1+cos2thetha=tanthetha
-----------
2sin(t)cos(t)/(1 + cos^2(t) - sin^2(t)) = tan(t)
2sin(t)cos(t)/(2cos^2(t)) = tan(t)
sin(t)cos(t)/cos^2(t) = tan(t)
sin(t)/cos(t) = tan(t)
I'm convinced.